OpenStudy (anonymous):

x • x = ? My answer is 2x Am I right? The other options are x^2 x(x + 2) x + 2

OpenStudy (baldymcgee6):

x*x = x^2 x + x = 2x

OpenStudy (anonymous):

6*6 = 36 = 6^2

OpenStudy (anonymous):

Thanks guys! :)